OpenStudy (anonymous):

Factor Completely: 36x2 - 25

OpenStudy (anonymous):

(6x + 5)(6x - 5) or (6x - 5)(x + 5)

OpenStudy (lgbasallote):

difference of two squares.. \(\large (\sqrt{36x^2} - \sqrt{25})(\sqrt{36x^2} + \sqrt{25})\)